Please help me find the equation in which the line has a X-intercept of 5 and a slope of 3
Answer by Alan3354(69443)

Use y = mx + b where m is the slope and b is the y-intercept, and the point.
m = 3 and the point (5,0)
Use y = mx + b and the point to find b
0 = 3*5 + b
b = -15
--> y = 3x - 15